OpenStudy (anonymous):

The height of a triangle is 4 m more than twice the length of the base. The area of the triangle is 63 m2. Find the height of the triangle.

OpenStudy (espex):

If length b, h=4+2b, A=1/2bh = 63 \[^{1}_{2}bh=63\] \[^{1}_{2}\times b \times(4+2b)=63\] \[b \times (4+2b)=126\] \[4b+2b ^{2}=126\] Divide both sides by 2 \[b ^{2}+2b=63\] \[b ^{2}+2b -63=0\] Plug into the quadratic equation and solve for b b=-9, b=7 Check: 1/2 7*(4+2*7)=63 1/2*7(4+14)=63 1/2*126=63 63=63